Grade 12th passMechanics

During an adiabatic process the volume of a gas is found to be proportional to square root of its absolute temperature. The ratio of specific heat capacity is...?

Abdul Salam
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
P^(1-gamma)*T^(gamma)=constant,
=> P^(gamma-1) = k* T^(gamma),
=> P = k' * T^(gamma/[gamma-1])
Given P= k' * T^2
Therefore,
 gamma/(gamma-1)=2,
=> gamma=2,
=> Cp/Cv=2